OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

n + n + 1 + n + 2 = 267. 3n + 3 = 267 n = ?

OpenStudy (anonymous):

Largest will be n+2.

OpenStudy (anonymous):

x = 88, 88 + 2 = 90 its 90 :)

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

oh okat haha thanks!

OpenStudy (anonymous):

yw :)